Find the critical values for a sample with n = 6 and α = 0.01 if H1 : μ <20.

Group of answer choices

A.-3.365

B.-2.977

C.-3.625

D.-3.761
Answer #1

solution:

Degrees of freedom = df = n - 1 = 6 - 1 = 5

=0.01

t ,df = t0.01,5 = -3.365 ( using student t table)